# Generating Reservoir State Descriptions with Random Matrices

### v1

@@ Abstract
We demonstrate a novel approach to reservoir computer measurements through the use of a simple quantum system and random matrices to motivate how atomic-scale devices might be used for real-world computing applications. In our approach, random matrices are used to construct reservoir measurements, introducing a simple, scalable means for producing state descriptions. In our studies, systems as simple as a five-atom Heisenberg spin chain are used to perform several tasks, including time series prediction and data interpolation. The performance of the measurement technique as well as their current limitations are discussed in detail alongside an exploration of the diversity of measurements yielded by the random matrices. Additionally, we explore the role of the parameters of the spin chain, adjusting coupling strength and the measurement dimension, yielding insights into how these learning machines might be automatically tuned for different problems. This research highlights the use of random matrices as measurements of simple quantum systems for natural learning devices and outlines a path forward for improving their performance and experimental realisation.
Keywords Random matrix theory, spin chain, reservoir computing, extreme learning machines
@@ 1 Introduction
Computing with physical systems such as particle swarms [1, 2] or lasers [3, 4] has emerged as a promising candidate for a new wave of technologies built around better utilizing the world around us. These approaches to computation, broadly captured under the term Extreme Learning Machines (ELMs) [5], of which Reservoir Computing (RC) can be seen as a subset, suggest that the majority function of highly parameterized learning devices is simply the projection of information into high dimensions, from there, a much smaller, trainable readout layer may be tuned to make predictions on this projection. One promising research direction is the use of quantum devices such as spin-lattices [6], superconducting qubits [7, 8], or artificial atoms [9], which are viable candidates for future, real-world devices due to their size and solid-state nature. A convenient aspect of utilising quantum devices for this form of computing comes in their power-law scaling of states as a function of particles. This factor makes quantum devices well-suited for projecting input signals into high-dimensional spaces, which can then be read off for predictions. To this end, there have been numerous demonstrations of the efficacy of quantum RC methods ranging from excited state energy predictions in molecules [10], quantum reinforcement learning [11] and object or image classification [7, 12] to mention just a few. A critical problem, however, is the construction of the state descriptions of the learning machine. This description must comprise a large set of numbers that provide a complete or over-complete representation of the system. In the case of quantum systems, the space of well-defined measurements that could be used in this description is small, and in most cases, only the Pauli spin matrices are used on individual spins [13, 14, 15] to construct such a state description. This approach results in small state representations and, therefore, less expressive measurements. In this work, we suggest utilising random measurements on the reservoir as an alternative to the Pauli spin matrices. This approach is motivated by the fact that the state description passed into the reservoir’s readout layer need not be interpretable; rather, it should be sensitive to changes in the internal state. Therefore, rather than restricting measurements to the Pauli spin matrices, we explore a fixed set of random measurements. This can be achieved by constructing random hermitian matrices, which can then act on the systems’ state in the presence of a driving field. In this way, large numbers of diverse measurements can be made on the reservoir, producing larger state descriptions, which, we argue, aids in better reservoir computing.
@@
Incorporating Random Matrix Theory (RMT) into the framework of quantum reservoir computing (QRC) here presents a novel approach to harnessing the intrinsic complexities of quantum systems for computational tasks. RMT, with its roots deeply embedded in statistical physics and mathematical theories, provides a versatile tool for understanding the statistical behaviours of complex systems [16, 17, 18, 19], particularly those governed by quantum mechanics. The application of RMT in this context is motivated by the unique ability of random matrices to capture the dynamic properties of quantum systems, thereby enriching the state descriptions used in reservoir computing. Using random matrices to generate reservoir state descriptions in quantum systems can increase the richness of a reservoir’s state space can significantly enhance its computational capacity. This strategy not only expands the expressive power of the state representations but also aligns with the core philosophy of reservoir computing, where the non-linearity and high-dimensionality of the reservoir’s dynamics are crucial for processing and predicting complex time-series data.
@@
The paper is structured as follows. We start with a brief overview of the state of the art in the field of quantum reservoir computing and extreme learning regimes. Afterwards, the theory associated with our approach is introduced and explained. We then present the results of our investigations on several test problems which aim to both demonstrate the capability of the method but also better understand the effects of tuning certain model parameters. Finally, we discuss the limitations and outlook of the methods before concluding.
@@ 1.1 State of the Art
Quantum reservoir computing and unconventional computing with quantum devices as seen a rise in interest in the preceding years [20, 21, 22]. The possibility of computing on such small scales and finding near-term benefits from quantum computing systems has driven scientists to make remarkable strides in the fields. Currently, most effort is placed on identifying the best mediums for this kind of computing at a quantum scale. Efforts have included the use of superconducting qubits [7, 8] in an attempt to adjust the application of quantum computers to reservoir computing or vastly reducing the size of quantum circuits [23], spin lattices [6] and crystals via spin-waves [24] and in arrays of Rydberg atoms [9]. In the direction of optical approaches, scientists have utilised quantum dot micropillar arrays [25], photonic memristors [12], closed-loop optical pulses [26], and photonic-based quantum recurrent neural networks [4]. Another research direction is better understanding the role of the Hilbert space in the efficacy of quantum reservoir computing and how one can maximally span the available state space. In this field, scientists have explored how the scaling of the Hilbert space can improve reservoir performance [27] as well as the role of certain kinds of quantum interactions such as entanglement [28]. Some work has looked directly at the measurement problem as in Mujal et. al. [29] where they study to use of intermittent measurements to produce a more reliable workflow for on-line RC. In their 2022 work, Domingo et. al [10] demonstrated that random quantum circuits could be used to construct state descriptions from molecular input data. Such an approach is similar in nature to the work presented here, albeit requiring a larger number of implemented quantum circuits. In all cases, these investigations have utilised the standard set of observables for the state description, typically the Pauli spin matrices. While this allows for a more interpretable state description, the important aspect of the representation is its sensitivity to changes with respect to the evolution of the reservoir. In other recent work by Haug et. al. [30] and Elben et. al. [31], random measurements were used to describe quantum support vector machines and quantum many-body systems, respectively, to great effect in a similar vein to what is presented here. This work introduces a more generalizable approach to state measurement wherein random matrices are used as measurements on the reservoir to construct a vector representation.
@@ 2.3 Neural Layer
where $i$ is the output, $x_{j}$ is the $j^{\text{th}}$ dimension of the signal and $w_{ij}$ is a prefactor connecting the two. In this work, we often found slightly improved results by allowing one additional layer with a ReLU non-linearity, still vastly reducing the size of the networks but increasing the complexity of the reservoir readout.
@@ 2.4 XXX Heisenberg Model
In this study, a five-atom spin chain is used as a physical reservoir, simulated in the QuTIP engine [36, 37]. The spin chain is modelled by the time-dependent Heisenberg Hamiltonian
@@
where $J$ is the spin coupling constant, $\sigma$ is an atom in the chain, $\langle ij\rangle$ denotes summation over spin pairs, and $h(t)$ is a time varying magnetic field. The driving signal is encoded into the reservoir through the magnetic field as it couples to the spins in the chain. Time evolution is simulated by solving the Lindbladian master equation [38, 39] implemented in QuTIP.
@@ 3.1 Observable Study
In all studies presented here, we are using systems of spins coupled with their nearest neighbours. In such composite systems, the possible states attainable scales by the power of the number of elements, i.e, for a combined state, $|\psi\rangle$
@@ 3.2 Cosine Wave
Figure 3: Scan over reservoir parameters for the prediction of the Cosine wave. $\tau$ is a period of the cosine wave and the colour corresponds to the Pearson correlation coefficient between the reservoir predictions and the ground truth data. Red indicates a better correlation, whereas blue indicates anti-correlation.
Looking first at the prediction length, we see that the performance decreases as the reservoir must predict further into the future. This is expected. However, it is interesting to note that the model can make relatively stable predictions even at 2.5 times the period. The relationship between the reservoir parameters is perhaps of more interest. We see a stable region of coupling strengths and state dimensions within which the reservoir can produce well-correlated predictions to the underlying signal. This region is relatively unchanged with respect to prediction length. Interestingly, there appears to be a slight tradeoff between coupling strength and state dimension for the higher spin couplings. As this exists for all prediction lengths, it does appear that a larger state dimension may be required to map the more strongly correlated reservoirs correctly. It should be noted that these coupling strengths do not explore any strongly correlated regime, as they respond only to mild driving signals. This is largely due to limitations on the simulation time but is an avenue that should be explored in future work.
@@ 3.3 Stock Data Interpolation
Figure 4: (a) Interpolation of the Dow30 price using the qRC and a 70 % split. (b) Pearson correlation coefficients for qRC, Cubic spline, and Cubic Hermite spline methods as a function of the train fraction. (c) Sample Correlation diagrams from the qRC fitting on 70 % train splits for test (left) and train (right) data.
@@
Interestingly, the qRC method can perform excellent interpolation on as little as 20 % of the data. This is comparable to the other spline methods, although these methods do not use a high-dimensional representation of the reservoir but rather simple 1-D interpolation. We also attempted to perform direct future prediction on the input data; however, the approach appears to be much more capable of interpolation than extrapolation. This also suggests that some memory elements are poorly characterised in the current implementation and could be improved in future work.
3.4 Mackey-Glass Function
The final test we ran used the spin chain to predict the future dynamics of the Mackey-Glass time series [40] with parameters $\{\beta,\gamma,\tau,n\}=\{0.2,0.1,17,10\}$. The time series is generated using a fourth-order Runge-Kutta integrator [41, 42] at a $0.1s$ time step. In order to reduce the simulation time, the resultant time series was sampled every 20 time steps and these samples passed into the reservoir. In this study, the function was imposed onto the reservoir through the magnetic field with a coupling of $0.1\cdot\pi$ and using a state dimension of $1000$. We perform a single-step (2 s in real-time), open-loop prediction on the driving function.
Figure 5: Open-loop single-step (2 s) prediction on test and train data for the Mackey-Glass Function
Figure 5 shows the train and test data in the Mackey-Glass function prediction. While the reservoir can learn some dynamics after the training domain, more work is needed to extract better generalisations and cleaner predictions. What is clear, however, is that the random matrices could capture sufficient information to perform predictions on the non-linear time series.
@@ 4 Practical Realisation
The results above show a brief demonstration of what is possible with this approach to reservoir measurement. The question remains on the applicability of such a method to real hardware. With the current means of quantum device measurement, performing expectations over hundreds of measurements is not feasible. However, there are approaches to overcoming this problem. In many real devices, scientists rely on measurements being performed further away from the driving mechanism such that many can be performed without disturbing the main dynamics. In these cases, the driving signal is oftentimes still repeated such that the required statistics can be achieved, resulting in a workflow that is, again, not real-time. However, much as we today utilise chips limited by a bit architecture, it can be imagined that such a procedure might extend to the reservoir computing world where small devices, stacks of spin lattices for example, are designed such that hundreds of measurements over thousands of ensembles can be performed simultaneously
@@ 5 Conclusion and Outlook
We have introduced a new approach to measuring quantum reservoir computers through the use of random matrices. Our method can be scaled to arbitrary state dimensions and shows good performance in the test cases. While the results shown here correspond to quantum systems, any reservoir computer paradigm could be substituted so long as the measurement problem can be phrased in terms of matrix operations. As a first investigation, we showed the spectrum and observable range of the random matrices on random system states and found that partial traces over the full system may improve the diversity of the observables that can be used in reservoir computing. We then demonstrated the efficacy of our method on the problem of time-series prediction for a sine wave and the Mackey-Glass time series, as well as data interpolation on historical Dow30 data. Our results suggest that the use of random matrices as a means of state description works well in the tested cases. However, the reservoirs implemented for the investigation were not always suitable for more complex problems. One avenue for future research is the study of much more strongly correlated spin systems or alternative reservoirs with a better learning capacity. Furthermore, larger systems should be probed to better understand the influence of correlations spanning more degrees of freedom, possibly also in experiments.
Acknowledgements
C.H and S.T acknowledge financial support from the German Funding Agency (Deutsche Forschungsgemeinschaft DFG) under Germany’s Excellence Strategy EXC 2075-390740016. C.H and S.T acknowledge financial support from the German Funding Agency (Deutsche Forschungsgemeinschaft DFG) under the Priority Program SPP 2363, “Utilization and Development of Machine Learning for Molecular Applications - Molecular Machine Learning” Project No. 497249646. C.H and S.T acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG. C.H and S.T acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Compute Cluster grant no. 492175459.

### v2

@@ Abstract
We demonstrate a novel approach to reservoir computer measurements using random matrices. We do so to motivate how atomic-scale devices might be used for real-world computing applications. Our approach uses random matrices to construct reservoir measurements, introducing a simple, scalable means for producing state descriptions. In our studies, two reservoirs, a five-atom Heisenberg spin chain, and a five-qubit quantum circuit, perform time series prediction and data interpolation. The performance of the measurement technique and current limitations are discussed in detail alongside an exploration of the diversity of measurements yielded by the random matrices. Additionally, we explore the role of the parameters of the reservoirs, adjusting coupling strength and the measurement dimension, yielding insights into how these learning machines might be automatically tuned for different problems. This research highlights using random matrices to measure simple quantum reservoirs for natural learning devices and outlines a path forward for improving their performance and experimental realization.
@@ 1 Introduction
Computing with physical systems such as particle swarms [1, 2] or lasers [3, 4] has emerged as a promising candidate for a new wave of technologies built around better utilizing natural systems as learning devices. These approaches to computation, broadly captured under the term Extreme Learning Machines (ELMs) [5], of which Reservoir Computing (RC) can be seen as a subset, suggest that the majority function of highly parameterized learning devices is simply the projection of information into high dimensions, from there, a much smaller, trainable readout layer may be tuned to make predictions on this projection. One promising research direction is the use of quantum devices such as spin-lattices [6], superconducting qubits [7, 8], or artificial atoms [9], which are viable candidates for future, real-world devices due to their size and solid-state nature. A convenient aspect of utilizing quantum devices for this form of computing comes in their power-law scaling of states as a function of particles. This factor makes relatively small quantum devices well-suited for projecting input signals into high-dimensional spaces, which can then be read off for predictions. To this end, there have been numerous demonstrations of the efficacy of quantum RC methods ranging from excited state energy predictions in molecules [10], quantum reinforcement learning [11], and object or image classification [7, 12] to mention just a few. A critical problem, however, is the construction of the state descriptions of the learning machine. This description must comprise a large set of numbers that provide a complete or over-complete representation of the system. In the case of quantum systems, the space of well-defined measurements that could be used in this description is small, and in most cases, only the Pauli spin matrices are used on individual spins [13, 14, 15] to construct such a state description. This approach results in small state representations and, therefore, less expressive measurements. In this work, we suggest utilizing random measurements on the reservoir as an alternative to the Pauli spin matrices. This approach is motivated by the fact that the state description passed into the reservoir’s readout layer need not be interpretable; rather, it should be sensitive to changes in the internal state. Therefore, rather than restricting measurements to the Pauli spin matrices, we explore a fixed set of random measurements. This can be achieved by constructing random hermitian matrices, which can then act on the systems’ state in the presence of a time-varying input. In this way, large numbers of diverse measurements can be made on the reservoir, producing larger state descriptions, which, we argue, aids in better reservoir computing.
@@
Incorporating Random Matrix Theory (RMT) into the framework of quantum reservoir computing (QRC) here presents a novel approach to harnessing the intrinsic complexities of quantum systems for computational tasks. RMT, with its roots deeply embedded in statistical physics and mathematical theories, provides a versatile tool for understanding the statistical behaviors of complex systems [16, 17, 18, 19], particularly those governed by quantum mechanics. The application of RMT in this context is motivated by the unique ability of random matrices to capture the dynamic properties of quantum systems, thereby enriching the state descriptions used in reservoir computing. Using random matrices to generate reservoir state descriptions in quantum systems can increase the richness of a reservoir’s state space can significantly enhance its computational capacity. This strategy not only expands the expressive power of the state representations but also aligns with the core philosophy of reservoir computing, where the non-linearity and high-dimensionality of the reservoir’s dynamics are crucial for processing and predicting complex time-series data.
@@
The paper is structured as follows. We start with a brief overview of the state of the art in the field of quantum reservoir computing and extreme learning regimes. Afterward, the theory associated with our approach is introduced and explained. We then present the results of our investigations on several test problems which aim to both demonstrate the capability of the method but also better understand the effects of tuning certain model parameters. Finally, we discuss the limitations and outlook of the methods before concluding.
@@ 1.1 State of the Art
Quantum reservoir computing and unconventional computing with quantum devices as seen a rise in interest in the preceding years [20, 21, 22]. The possibility of computing on such small scales and finding near-term benefits from quantum computing systems has driven scientists to make remarkable strides in the fields. Currently, most effort is placed on identifying the best mediums for this kind of computing at a quantum scale. Efforts have included the use of superconducting qubits [7, 8] in an attempt to adjust the application of quantum computers to reservoir computing or vastly reduce the size of quantum circuits [23], spin lattices [6] and crystals via spin-waves [24] and in arrays of Rydberg atoms [9]. In the direction of optical approaches, scientists have utilized quantum dot micropillar arrays [25], photonic memristors [12], closed-loop optical pulses [26], and photonic-based quantum recurrent neural networks [4]. Another research direction is better understanding the role of the Hilbert space in the efficacy of quantum reservoir computing and how one can maximally span the available state space. In this field, scientists have explored how the scaling of the Hilbert space can improve reservoir performance [27] as well as the role of certain kinds of quantum interactions such as entanglement [28]. Some work has looked directly at the measurement problem as in Mujal et. al. [29] where they study to use of intermittent measurements to produce a more reliable workflow for online RC. In their 2022 work, Domingo et. al [10] demonstrated that random quantum circuits could be used to construct state descriptions from molecular-input data. Such an approach is similar to the work presented here, albeit requiring a larger number of implemented quantum circuits. In all cases, these investigations have utilized the standard set of observables for the state description, typically the Pauli spin matrices. While this allows for a more interpretable state description, the important aspect of the representation is its sensitivity to changes with respect to the evolution of the reservoir. In other recent work by Haug et. al. [30] and Elben et. al. [31], random measurements were used to describe quantum support vector machines and quantum many-body systems, respectively, to great effect in a similar vein to what is presented here. This work introduces a more generalizable approach to state measurement wherein random matrices are used as measurements on the reservoir to construct a vector representation.
@@ 2.3 Neural Layer
where $i$ is the output, $x_{j}$ is the $j^{\text{th}}$ dimension of the signal and $w_{ij}$ is a prefactor connecting the two. In this work, we often found slightly improved results by allowing one additional layer with a ReLU non-linearity, still vastly reducing the size of the networks but increasing the complexity of the reservoir readout. In the case of the spin chains, a single ReLU activation is used in the readout layer whereas in the quantum circuit reservoirs, no non-linearity was required.
@@ 2.4 XXX Heisenberg Model
In this study, the simplest reservoir used is a five-atom spin chain, simulated in the QuTIP engine [36, 37]. The spin chain is modeled by the time-dependent Heisenberg Hamiltonian
@@
where $J$ is the spin coupling constant, $\sigma$ is an atom in the chain, $\langle ij\rangle$ denotes summation over spin pairs, and $h(t)$ is a time varying magnetic field. The driving signal is encoded into the reservoir through the magnetic field as it couples to the spins in the chain. Time evolution is simulated by solving the Lindbladian master equation [38, 39].
2.5 Quantum circuit
Additionally, we utilize a quantum circuit of five qubits as a quantum reservoir. We sequentially encode sequences of data-points from a time series into the quantum state by angle-encoding [40]. In this method, a rotation gate is acting on qubits with an angle corresponding to the value of a data point. A further unitary operation on the quantum state is applied in order to realize a versatile state representation. The circuit is simulated using the Python library Pennylane [41].
@@ 3.1 Observable Study
In the studies presented here, we employ either coupled spins or qubits within a quantum circuit. In such composite systems, the possible states attainable scales by the power of the number of elements, i.e, for a combined state, $|\psi\rangle$
@@ 3.2 Cosine Wave
Figure 3: Scan over reservoir parameters for the prediction of the Cosine wave. $\tau$ is a period of the cosine wave and the color corresponds to the Pearson correlation coefficient between the reservoir predictions and the ground truth data. Red indicates a better correlation, whereas blue indicates anti-correlation.
Looking first at the prediction length, we see that the performance decreases as the reservoir must predict further into the future. This is expected. However, it is interesting to note that the model can make relatively stable predictions even at 2.5 times the period. The relationship between the reservoir parameters is perhaps of more interest. We see a stable region of coupling strengths and state dimensions within which the reservoir can produce well-correlated predictions to the underlying signal. This region is relatively unaffected by prediction length. Interestingly, there appears to be a slight tradeoff between coupling strength and state dimension for the higher spin couplings. As this exists for all prediction lengths, it does appear that a larger state dimension may be required to map the more strongly correlated reservoirs correctly. It should be noted that these coupling strengths do not explore any strongly correlated regime, as they respond only to mild driving signals. This is largely due to limitations on the simulation time but is an avenue that should be explored in future work.
@@ 3.3 Stock Data Interpolation
Figure 4: (left) Interpolation of the Dow30 price using the qRC and a 70 % split. (top right) Pearson correlation coefficients for qRC, Cubic spline, and Cubic Hermite spline methods as a function of the train fraction. (bottom right) Sample Correlation diagrams from the qRC fitting on 70 % train splits for test (left) and train (right) data.
@@
Interestingly, the qRC method can perform excellent interpolation on as little as 20 % of the data. This is comparable to the other spline methods, although these methods do not use a high-dimensional representation of the reservoir but rather simple 1-D interpolation. We also attempted to perform direct future prediction on the input data; however, the approach appears to be much more capable of interpolation than extrapolation. This also suggests that some memory elements are poorly characterized in the current implementation and could be improved in future work.
3.4 Quantum Circuit Reservoir Computing on the Mackey-Glass Series
For our final investigation, we examine a quantum circuit comprising 5 qubits as a reservoir. We train the reservoir to predict the future dynamics of the Mackey-Glass time series $P(t)$ [42] with parameters $\{\beta,\gamma,\tau,n\}=\{0.2,0.1,17,10\}$ and $P(0)=1.2$. The time series is generated using a fourth-order Runge-Kutta integrator [43, 44] at a $1s$ time step. The data is split into sequences of length $l$ and sequentially encoded into a state representation by the circuit depicted in Figure 5 (a). To realize a versatile state representation, parameter-dependent single and multi-qubit gates act on the quantum state. The parameters $\theta_{i}$ are sampled randomly from a uniform distribution and are fixed in the subsequent process. Following the circuit’s execution, a batch of random Hermitian matrices’ expectation values is measured. These results are then fed into a classical linear layer, which is trained to predict a future data point. We conducted a parameter scan for different sequence lengths $l=\{4,8,12,16\}$, different prediction steps $\{1,10,50,1000\}$ and different numbers of measurements equally spaced on a logarithmic scale in the interval $[1,1000]$. We split the data set into 67% training and 33% test data and compute the Pearson Correlation between the actual and predicted test data of the trained reservoir. The best result of the parameter scan is shown in Figure 5 (b). The values on the test set are predicted by feeding a sequence of the test data into the reservoir. The model achieves a correlation of up to $0.999$. Figure 5 (c) displays the results of the parameter scan. It’s noticeable that with an increased number of measurements, the model attains a higher correlation on the test data. Two effects can contribute to this effect. First, the size of the classical output layer increases and the model may gain more flexibility in learning the data. Second, the model obtains access to a more comprehensive description of the state representation. Moreover, it becomes apparent that by extending the sequence length, the model’s performance appears to diminish for a comparable number of measurements. This suggests the necessity for a more expressive model architecture to effectively capture longer sequences. Lastly, we notice that even for large prediction lengths of 50 or 100 steps, the model can perform predictions on the non-linear time series with a high accuracy.
Figure 5: (a) Quantum circuit used in this study. A sequence of data points is sequentially encoded into the quantum state of the five-qubit system. Single and multi-qubit gates with fixed rotation parameters $\theta_{i}$ act on the quantum state. At the end of the circuit, the expectation values of random Hermitian matrices are measured and fed into a linear layer. (b) Mackey-glass Function and the predictions of the trained model with the highest correlation score on the test data. (c) Result of the parameter scan for different sequence lengths, different prediction lengths, and different numbers of random hermitian measurements. Shown is the correlation on the test data for the trained models.
@@ 4 Practical Realisation
The results above demonstrate the potential of this reservoir measurement approach. However, its practical application to real hardware remains a question. Currently, performing expectations over hundreds of measurements using quantum devices is not feasible. Nonetheless, there are ways to overcome this challenge. In many real devices, scientists conduct measurements further away from the driving mechanism, allowing multiple measurements to be performed without disturbing the primary dynamics. In these cases, the driving signal is often repeated to obtain the required statistics, resulting in a non-real-time workflow. Looking to the future, just as we utilize chips limited by a bit architecture, it’s intriguing to imagine that this might extend to reservoir computing, where small devices like stacks of spin lattices are designed to perform hundreds of measurements over thousands of ensembles simultaneously.
@@ 5 Conclusion and Outlook
We have introduced a new approach to measuring quantum reservoir computers using random matrices. Our method can be scaled to arbitrary state dimensions and performs well in test cases utilizing two quantum reservoirs, a Heisenberg spin chain, and a quantum circuit. While the results shown here correspond to quantum systems, any reservoir computer paradigm could be substituted as long as the measurement problem can be phrased in matrix operations. As a first investigation, we showed the spectrum and observable range of the random matrices on random system states. We found that partial traces over the entire system may improve the diversity of the observables that can be used in reservoir computing. We then demonstrated the efficacy of our method on the problem of time-series prediction for a sine wave and the Mackey-Glass time series, as well as data interpolation on historical Dow30 data. Our results suggest that random matrices work well for different reservoirs for state description. One avenue for future research is the study of much more strongly correlated spin systems or alternative reservoirs with a better learning capacity. Furthermore, larger reservoirs should be probed to better understand the influence of correlations spanning more degrees of freedom, possibly also in experiments.